# Examining the influence of anisotropy on the fundamental mode of nonradial oscillation in quark stars on a complete general relativistic scheme

### v1

@@ Appendix B Perturbation functions near $r=0$
$\hskip-11.38092pt{\tilde{X}}(0)=(\rho_{{}_{0}}+p_{r{{}_{0}}})e^{\Psi_{0}}\left[-\frac{\omega^{2}e^{-2\Psi_{{}_{0}}}}{\ell}{\tilde{W}}(0)+\frac{4\pi}{3}\left(\rho_{{}_{0}}+3p_{r_{0}}\right){\tilde{W}}(0)-\frac{{\tilde{K}}(0)}{2}\right]+\alpha p_{r_{0}}\left[{\tilde{K}}(0)\right$.
$\left.-\frac{16\pi}{3}\rho_{{}_{0}}{\tilde{W}}(0)\right]e^{\Psi_{0}},$ (B.6)
@@
$\hskip-11.38092pt\rho_{{}_{2}}=-\frac{(\rho_{{}_{0}}+p_{r_{0}})^{2}}{2\Gamma\,p_{r_{0}}}\Psi_{{}_{2}},$ (B.13)
$\hskip-11.38092ptp_{r_{2}}=-\frac{4\pi}{3}(\rho_{{}_{0}}+3p_{r_{0}})(\rho_{{}_{0}}+p_{r_{0}})+\frac{16\pi}{3}\alpha\rho_{{}_{0}}\,p_{r_{0}},$ (B.14)
$\hskip-11.38092pt\Psi_{{}_{2}}=\frac{8\pi}{3}(\rho_{{}_{0}}+3p_{r_{0}})-\frac{16\pi\alpha}{3}\frac{\rho_{{}_{0}}p_{r_{0}}}{\rho_{{}_{0}}+p_{r_{0}}},$ (B.15)
$\hskip-11.38092pt\sigma_{{}_{2}}=\frac{16\pi\alpha}{3}\rho_{{}_{0}}\,p_{r_{0}},$ (B.16)
$\hskip-11.38092ptp_{r_{4}}=-\frac{2\pi}{5}(\rho_{{}_{2}}+5p_{r_{2}})(\rho_{0}+p_{r_{0}})-\frac{2\pi}{3}\left(\rho_{{}_{0}}+3p_{r_{0}})(\rho_{{}_{2}}+p_{r_{2}}\right)-\frac{32\pi^{2}}{9}\rho_{{}_{0}}(\rho_{{}_{0}}+3p_{r_{0}})(\rho_{{}_{0}}+p_{r_{0}})$
$\hskip-11.38092pt+8\pi\alpha\left(\frac{\rho_{{}_{2}}p_{r_{0}}}{5}+\frac{\rho_{{}_{0}}p_{r_{2}}}{3}\right),$ (B.17)
$\hskip-11.38092pt\Psi_{{}_{4}}=\frac{1}{(\rho_{{}_{0}}+p_{r_{0}})}\left[8\pi\alpha\left(\frac{\rho_{{}_{2}}p_{r_{0}}}{5}+\frac{\rho_{{}_{0}}p_{r_{2}}}{3}\right)-\frac{\Psi_{{}_{2}}}{2}(\rho_{{}_{2}}+p_{r_{2}})-2p_{r_{4}}\right],$ (B.18)
$\hskip-11.38092pt\sigma_{{}_{4}}=\frac{16\pi\alpha}{3}\rho_{{}_{0}}p_{r_{2}}.$ (B.19)
Note that all coefficients shown in Eqs. (B.13)-(B.19) are concord with those obtained for the isotropic case, since when $\alpha=0$, Eqs. (B.14)-(B.15) and (B.17)-(B.18) reproduce the equalities reported in Ref. [19] and with the equalities (B.16) and (B.19) being identically null. Moreover, by evaluating the perturbation equation to the second-order terms $\left[{\tilde{H}}_{1}^{\prime\prime}(0),{\tilde{K}}^{\prime\prime}(0),{\tilde{W}}^{\prime\prime}(0),{\tilde{X}}^{\prime\prime}(0)\right]$, we find that the following relationships must hold:
@@
$\hskip-11.38092pt+8\pi(p_{r_{0}}+\rho_{{}_{0}})Q_{1}+\frac{1}{2}Q_{0}-\frac{8\pi}{\ell}\left(p_{r_{2}}+\rho_{{}_{2}}+\frac{16\pi\alpha}{3}\rho_{{}_{0}}p_{r_{0}}\right){\tilde{W}}(0),$ (B.20)
@@
$\hskip-11.38092pt\frac{\ell+2}{2}{\tilde{X}}^{\prime\prime}(0)+e^{\Psi_{0}}(p_{r_{0}}+\rho_{{}_{0}})\left[\frac{\omega^{2}}{2}e^{-2\Psi_{{}_{0}}}+2\pi\left(p_{r_{0}}+\rho_{{}_{0}}\right)-\frac{\ell+2}{2}\Psi_{{}_{2}}+\frac{\ell+1}{2}\frac{\sigma_{{}_{2}}}{p_{r_{0}}+\rho_{{}_{0}}}\right]{\tilde{W}}^{\prime\prime}(0)$
$\hskip-11.38092pt+\alpha e^{\Psi_{{}_{0}}}p_{r_{0}}{\tilde{K}}^{\prime\prime}(0)+\frac{\ell(\ell+1)}{8}e^{\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}}){\tilde{H}}_{1}^{\prime\prime}=\left[\frac{\ell}{2}\left[(p_{r_{0}}+\rho_{{}_{0}})\Psi_{{}_{2}}+p_{r_{2}}+\rho_{{}_{2}}+\sigma_{{}_{2}}\right]\right$.
$\hskip-11.38092pt\left.+\frac{d\rho}{dp_{r}}\sigma_{{}_{2}}\right]\frac{{\tilde{X}}(0)}{p_{r_{0}}+\rho_{{}_{0}}}+e^{\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})\biggl{[}\left(-\Psi_{{}_{2}}+\frac{3\sigma_{{}_{2}}}{p_{r_{0}}+\rho_{{}_{0}}}\right){\tilde{K}}(0)-\left(\frac{\alpha p_{r_{0}}}{p_{r_{0}}+\rho_{{}_{0}}}+\frac{1}{4}\right)Q_{0}$
$\hskip-11.38092pt-\frac{\omega^{2}}{2}e^{-2\Psi_{{}_{0}}}{\tilde{H}}_{1}(0)-\frac{\ell(\ell+1)}{2}\left(\Psi_{{}_{2}}-\frac{\sigma_{{}_{2}}}{p_{r_{0}}+\rho_{{}_{0}}}\right)Q_{1}+\biggl{[}(\ell+2)\Psi_{{}_{4}}+\frac{\ell\sigma_{{}_{4}}}{2(p_{r_{0}}+\rho_{{}_{0}})}$
$\hskip-11.38092pt-2\pi\left(p_{r_{2}}+\rho_{{}_{2}}+\sigma_{{}_{2}}\right)-\frac{16\pi^{2}}{3}\rho_{{}_{0}}(p_{r_{0}}+\rho_{{}_{0}})-\frac{4\pi}{3}\rho_{{}_{0}}\Psi_{{}_{2}}+\omega^{2}e^{-2\Psi_{{}_{0}}}\left(\Psi_{{}_{2}}-\frac{4\pi}{3}\rho_{{}_{0}}\right)\biggr{]}{\tilde{W}}(0)\biggr{]}$ (B.22)
@@
$+\frac{\alpha}{2}\left[p_{r_{2}}{\tilde{K}}(0)+p_{r_{0}}\left(Q_{0}+{\tilde{K}}^{\prime\prime}(0)\right)\right]=-\frac{1}{2}e^{-\Psi_{{}_{0}}}\Psi_{{}_{2}}{\tilde{X}}(0)+\frac{1}{4}(p_{r_{2}}+\rho_{{}_{2}}){\tilde{K}}(0)$
$+\frac{1}{4}\left(p_{r_{0}}+\rho_{{}_{0}}\right)Q_{0}+\frac{1}{2}\alpha p_{r_{0}}\Psi_{{}_{2}}{\tilde{K}}(0)+\left[\frac{\omega^{2}}{\ell}e^{-2\Psi_{{}_{0}}}\left[\frac{1}{2}(p_{r_{2}}+\rho_{{}_{2}}+\sigma_{{}_{2}})-(p_{r_{0}}+\rho_{{}_{0}})\Psi_{{}_{2}}\right]\right$.
$\hskip-11.38092pt\left.+p_{r_{4}}-\frac{4\pi}{3}\rho_{{}_{0}}p_{r_{2}}\right]{\tilde{W}}(0)-\frac{1}{2}\omega^{2}e^{-2\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})Q_{1},$ (B.23)
@@
$\hskip-11.38092pt\left.-\ell(\ell+1)\left(\frac{2\pi}{3}\rho_{{}_{0}}+2\pi p_{r_{0}}\right)\right]{\tilde{H}}_{1}{(0)}-\frac{128\pi^{2}}{3}\alpha\rho_{{}_{0}}p_{r_{0}}{\tilde{W}}(0)\biggr{]},$ (B.24)
@@
It is important to mention that all equations presented in this section agree with those presented in [40, 19] for the isotropic case.

### v2

@@ Appendix B Perturbation functions near $r=0$
$\hskip-11.38092pt{\tilde{X}}(0)=(\rho_{{}_{0}}+p_{r{{}_{0}}})e^{\Psi_{0}}\left[-\frac{\omega^{2}e^{-2\Psi_{{}_{0}}}}{\ell}{\tilde{W}}(0)+\frac{4\pi}{3}\left(\rho_{{}_{0}}+3p_{r_{0}}\right){\tilde{W}}(0)-\frac{{\tilde{K}}(0)}{2}\right],$ (B.6)
@@
$\hskip-11.38092pt\Psi_{{}_{2}}=\frac{4\pi}{3}(\rho_{{}_{0}}+3p_{r_{0}}),$ (B.13)
$\hskip-11.38092pt\rho_{{}_{2}}=-\frac{(\rho_{{}_{0}}+p_{r_{0}})^{2}}{\Gamma\,p_{r_{0}}}\Psi_{{}_{2}},$ (B.14)
$\hskip-11.38092pt\sigma_{{}_{2}}=0,$ (B.15)
$\hskip-11.38092ptp_{r_{2}}=-(\rho_{{}_{0}}+p_{r_{0}})\Psi_{2},$ (B.16)
$\hskip-11.38092pt\Psi_{{}_{4}}=\frac{2\pi}{5}(\rho_{{}_{2}}+5p_{r_{2}})+\frac{32\pi^{2}}{9}\rho_{{}_{0}}(\rho_{{}_{0}}+3p_{r_{0}}),$ (B.17)
$\hskip-11.38092pt\sigma_{{}_{4}}=\left(\frac{16\pi\rho_{{}_{0}}}{3}\right)^{2}p_{r_{0}}\alpha,$ (B.18)
$\hskip-11.38092ptp_{r_{4}}=-(p_{r_{0}}+\rho_{{}_{0}})\Psi_{{}_{4}}-\frac{1}{2}(p_{r_{2}}+\rho_{{}_{2}})\Psi_{{}_{2}}+\frac{1}{2}\sigma_{{}_{4}}.$ (B.19)
Note that all coefficients shown in Eqs. (B.13)-(B.19) are concord with those obtained for the isotropic case. Since when $\alpha=0$, Eqs. (B.13)-(B.17) and (B.19) reproduce the equalities reported in Ref. [76] and with the equalities (B.15) and (B.18) being identically null. Moreover, by evaluating the perturbation equation to the second-order terms $\left[{\tilde{H}}_{1}^{\prime\prime}(0),{\tilde{K}}^{\prime\prime}(0),{\tilde{W}}^{\prime\prime}(0),{\tilde{X}}^{\prime\prime}(0)\right]$, we find that the following relationships must hold:
@@
$\hskip-11.38092pt+8\pi(p_{r_{0}}+\rho_{{}_{0}})Q_{1}+\frac{1}{2}Q_{0}-\frac{8\pi}{\ell}\left(p_{r_{2}}+\rho_{{}_{2}}\right){\tilde{W}}(0),$ (B.20)
@@
$\hskip-11.38092pt\frac{\ell+2}{2}\tilde{X}^{\prime\prime}(0)+\frac{\ell(\ell+1)}{8}e^{\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})\tilde{H}_{1}^{\prime\prime}(0)+e^{\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})\left[\frac{\omega^{2}}{2}e^{-2\Psi_{{}_{0}}}+2\pi(p_{r_{0}}+\rho_{{}_{0}})\right$.
$\hskip-11.38092pt\left.-\frac{\ell+2}{2}\Psi_{{}_{2}}\right]\tilde{W}^{\prime\prime}(0)=\frac{\ell}{2}\left(\Psi_{{}_{2}}+\frac{p_{r_{2}}+\rho_{{}_{2}}}{p_{r_{0}}+\rho_{{}_{0}}}\right)\tilde{X}(0)+e^{\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})\biggl[-\frac{\ell(\ell+1)}{2}\Psi_{{}_{2}}Q_{1}-\frac{1}{4}Q_{0}$
$\hskip-11.38092pt-\left(\frac{\omega^{2}}{2}e^{-2\Psi_{{}_{0}}}\right)\tilde{H}_{1}(0)-\Psi_{{}_{2}}\tilde{K}(0)+\biggl\{(\ell+2)\Psi_{{}_{4}}-2\pi(p_{r_{2}}+\rho_{{}_{2}})-\frac{16\pi^{2}}{3}\rho_{{}_{0}}(p_{r_{0}}+\rho_{{}_{0}})-\frac{4\pi}{3}\rho_{{}_{0}}\Psi_{{}_{2}}$
$\hskip-11.38092pt+\omega^{2}e^{-2\Psi_{{}_{0}}}\left(\Psi_{{}_{2}}-\frac{4\pi}{3}\rho_{{}_{0}}\right)-\frac{\sigma_{4}}{p_{r_{0}}+\rho_{{}_{0}}}\biggr\}\tilde{W}(0)\biggr],$ (B.22)
@@
[…]
$\left.\left.-(p_{r_{0}}+\rho_{{}_{0}})\Psi_{{}_{2}}\right]+p_{r_{4}}-\frac{4\pi}{3}\rho_{{}_{0}}p_{r_{2}}\right]{\tilde{W}}(0)-\frac{1}{2}\omega^{2}e^{-2\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})Q_{1},$ (B.23)
@@
$\hskip-11.38092pt\left.-\ell(\ell+1)\left(\frac{2\pi}{3}\rho_{{}_{0}}+2\pi p_{r_{0}}\right)\right]{\tilde{H}}_{1}{(0)}\biggr],$ (B.24)
@@
It is important to mention that all equations presented in this section agree with those presented in [51, 76] for the isotropic case. In addition, the expansion of Eq. (B.21) to the leading order is the same as Eq. (B.6).